2(5-2y)-3(1-y)=y+1

Answer by tommyt3rd(5050)

2(5-2y)-3(1-y)=y+1
10-4y-3+3y=y+1
10-3-4y+3y=y+1
-y+y+7-1=y+y+1-1
2y=6
y=3